Solve the given linear system by any method. 3x1 + X2 + X3 + X4 = 0, 5x1 - X2 + X3 - X4 = 0


1
Expert's answer
2022-02-21T17:38:59-0500

"\begin{bmatrix}
 3& 1&1&1 \\
 5&-1&1&-1
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2\\x_3\\x_4
\end{bmatrix}=0"


"\frac{1}5R_2,\frac{1}3R_1"


"\begin{bmatrix}
 1& ⅓&⅓&⅓ \\
 1 & -\frac{1}5&\frac{1}5&-\frac{1}5
\end{bmatrix}"


"\frac{-15}8(R_2-R_1)"


"\begin{bmatrix}
 1& ⅓&⅓&⅓\\
 0&1 & ¼&1
\end{bmatrix}"


"R_1-\u2153R_2"


"\begin{bmatrix}
 1& 0&¼&0 \\
 0 & 1&¼&1
\end{bmatrix}"


"x_1+\u00bcx_3=0"


"x_2+\u00bcx_3+x_4=0"


"x_1=-\u00bcx_3"


"x_2=-\u00bcx_3-x_4"